OpenStudy (anonymous):

how to solve x^3(y'-x)=y^2

OpenStudy (anonymous):

whats that? y prime?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is this DE

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok did you arived yet in dy/dx -(1/x^3)y^2 =x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

dy/dx -(1/x^3)y^2 =x integral of dy=y=integral of [ x+(x^-3)(y^2)]dx try intgrating that

OpenStudy (anonymous):

y=integral of [ x+(x^-3)(y^2)]dx y=(x^4-y^2)/2x^2